Rita's tuition bill for last semester was $4360. If she paid $6052 in tuition this semester, what was the percentage increase on her tuition?
Answer by amarjeeth123(570)

Rita's tuition bill for last semester was $4360.
She paid $6052 in tuition this semester.
increase=6052-4360=1692
% increase=1692/4360*100=38.8%